OpenStudy (anonymous):

please help me One number is 4 more than another, and their sum is 60. What is the smaller number? 26 28 32

OpenStudy (anonymous):

\[\ Let \ smaller\ number \ be \ x\] \[\ then \ according \ problem, \ other \ number \ = x+4\] \[\ given \ thesum \ of \ number = 60\] \[\therefore \ x+(x+4) =60\] \[\implies 2x+4=60 \implies 2x=60-4=56\] \[\therefore x=\frac{ 56 }{ 2 }=28\] \[\therefore smaller \ number \ is \ 28\\ and \ other \ number\ i.e. (x+4) \ is \ 32\]

OpenStudy (anonymous):

thank u